# Qimax: Efficient quantum simulation via GPU-accelerated extended stabilizer formalism

### v1

@@ Appendix C Single-qubit Pauli measurement
Let $\mathbb{P}_{n,k,b}=\underbrace{I\otimes\cdots\otimes|b\rangle\langle b|_{k}\otimes\cdots\otimes I}_{n\;\text{terms}}=\frac{1}{2}((-1)^{b}Z_{n,k}+I^{\otimes n})$ for $k\in[n]$. When measuring the $k^{\text{th}}$ qubit of an $n$-qubit state $|\psi\rangle$ using projectors $\{\mathbb{P}_{n,k,0},\mathbb{P}_{n,k,1}\}$, we get two possible outcomes: 0 with probability $p_{k,0}$ and 1 with probability $p_{k,1}$. It follows that $p_{k,0}=\mathrm{Tr}(\mathbb{P}_{n,k,0}|\psi\rangle\langle\psi|)$ where $\mathrm{Tr}$ is the trace mapping. As shown in Eq. 10 and Eq. 11, the density operator $|\psi\rangle\langle\psi|$ can be written using generators i.e., $|\psi\rangle\langle\psi|=\frac{1}{2^{n}}\sum_{\mathbb{P}\in G}\mathbb{P}$. The probability $p_{k,0}$ can be obtained as:

### v2

@@ Appendix C Single-qubit Pauli measurement
Let $\mathbb{P}_{n,k,b}=\underbrace{I\otimes\cdots\otimes|b\rangle\langle b|_{k}\otimes\cdots\otimes I}_{n\;\text{terms}}=\frac{1}{2}((-1)^{b}Z_{n,k}+I^{\otimes n})$ for $k\in[n]$. When measuring the $k^{\text{th}}$ qubit of an $n$-qubit state $|\psi\rangle$ using projectors $\{\mathbb{P}_{n,k,0},\mathbb{P}_{n,k,1}\}$, we get two possible outcomes: 0 with probability $p_{k,0}$ and 1 with probability $p_{k,1}$. It follows that $p_{k,0}=\mathrm{Tr}(\mathbb{P}_{n,k,0}|\psi\rangle\langle\psi|)$ where $\mathrm{Tr}$ is the trace mapping. As shown in (10), the density operator $|\psi\rangle\langle\psi|$ can be written using generators i.e., $|\psi\rangle\langle\psi|=\frac{1}{2^{n}}\sum_{\mathbb{P}\in G}\mathbb{P}$. The probability $p_{k,0}$ can be obtained as: